OpenStudy (anonymous):

For an ellipse, where k is the sum of distance from any P(x,y) to its focal points, for what values of k will the locus be an ellipse, a line segment(what is the equation of the line), or the nullset?

OpenStudy (anonymous):

What about for hyperbolas, k=difference of distance between P and F F'. What values of k will the locus be a hyperbola,pair of rays(equation of each ray), null set? Can you have negative distance?

OpenStudy (compassionate):

Try asking this in MIT 8.01 or MIT Physics.

OpenStudy (anonymous):

I don't understand, the sum of the distances to the foci is equal to 2a ie Pf1 + Pf2 = 2a That's a definition for an ellipse.

OpenStudy (anonymous):

ok how about for what values of k, sum of distance between focal points and points(x,y), will the locus of the point be an ellipse, a line segment, or null set. e.g. let distance between foci=2c ellipse:k>2c line segment:k=2c y=0 F'(x)<=x<=F(x), x=0 F'(y)<=y<=F(x) nullset:k<2c is that right

OpenStudy (anonymous):

asdsadasd

OpenStudy (anonymous):

did i accidentally close my question..

OpenStudy (anonymous):

It's Ok, I was on the phone, just reading it now...

OpenStudy (anonymous):

"for what values of k, sum of distance between focal points and points(x,y), will the locus of the point be an ellipse, a line segment, or null set." The sum you are talking about is equal to 2a

OpenStudy (anonymous):

okso.. a^2=b^2+c^2 for ellipse so a>c.. yeah?

OpenStudy (anonymous):

and 2c>2a .: k<2c for hyperbola

OpenStudy (anonymous):

Usually it is written c^2 = a^2-b^2 but I prefer to think directly of the eccentricity e ie focus plusminus ae and e = sqrt(1-b^2/a^2)

OpenStudy (anonymous):

or 1+b^2/a^2 for hyperbola

OpenStudy (anonymous):

so.. for hyperbolas for what values of k will it exist as ahyperbola pair of ray or null set

OpenStudy (anonymous):

Define k

OpenStudy (anonymous):

PF'-PF=k

OpenStudy (anonymous):

P(x,y) F'(focal points)

OpenStudy (compassionate):

I should have never posted on this thread. I keep getting notifications. :(

OpenStudy (anonymous):

sorry :x

OpenStudy (anonymous):

hyperbola

OpenStudy (anonymous):

how about.. for what values of k is the locus of P(x,y) PF'-PF=k a hyperbola, pair of rays, null set.

OpenStudy (anonymous):

I think what you should do is post your question as a new question (this will stop the notifications and bring other people to the question)